OpenStudy (anonymous):

lim x / x->∞ sqrt x+1

OpenStudy (anonymous):

\[\frac{ x }{ \sqrt{x+1} }\]

geerky42 (geerky42):

Do you know English? No?

OpenStudy (anonymous):

so so

OpenStudy (anonymous):

do you can help me ??

geerky42 (geerky42):

Do you know which approaches infinity faster? \(x\) or \(\sqrt{x+1}\) Can you understand me?

OpenStudy (anonymous):

lim x->∞ \[\frac{ x }{ \sqrt{x+1} }\]

OpenStudy (anonymous):

understand now ??

geerky42 (geerky42):

@SithsAndGiggles

geerky42 (geerky42):

I understand. i try say something, but hard for us understand

OpenStudy (anonymous):

I explained to you how to do this problem last time you asked, but you replied with the same problem again, I'm not sure what exactly you need.

OpenStudy (anonymous):

\[\begin{align*}\lim_{x\to\infty}\frac{x}{\sqrt{x+1}}\cdot\frac{\sqrt{x+1}}{\sqrt{x+1}}&=\lim_{x\to\infty}\frac{x\sqrt{x+1}}{x+1}\\ &=\lim_{x\to\infty}\frac{\sqrt{x+1}}{1+\dfrac{1}{x}}\\ &\to\frac{\infty}{1+0}=\infty \end{align*}\]